# Polynomial-Time Near-Optimal Estimation over Certain Type-2 Convex Bodies

### v1

@@ 4 Robust Heavy-Tailed Extension
We are interested in estimating $\mu$ in a minimax optimal way. In order to derive the rate peng2025noisy, suppose that $N\gtrsim\sup_{\delta>0}\log M^{\operatorname{loc}}_{K}(\delta)$. Since $n\gtrsim\log M^{\operatorname{loc}}_{K}(\delta)$ this assumption is implied if $N\gtrsim n$. On the other hand, since we assume that the set $K$ is well-balanced, we have that $rB_{2}\subseteq K$ and therefore, when $\delta\lesssim r$, $\log M^{\operatorname{loc}}_{K}(\delta)\gtrsim n$. Thus, in this section, we will assume the assumption: $N\gtrsim n$. In that case, as shown by peng2025noisy the minimax rate $\inf_{\hat{\mu}}\sup_{\mu\in K}\sup_{\mathcal{C}}\mathbb{E}\|\hat{\mu}-\mu\|_{2}^{2}$ is $\asymp\eta^{\star 2}\wedge d^{2}$ where $\eta^{\star}=(\delta^{\star}\vee\sqrt{\varepsilon}\sigma)$ where $\delta^{\star}:=\sup\{\delta\geq 0:N\delta^{2}/\sigma^{2}\leq\log M^{\operatorname{loc}}_{K}(\delta)\}$, and where the inner most supremum over $\mathcal{C}$ indicates taking supremum over all possible adversarial schemes.
@@
We will modify several steps in our algorithm. Specifically, after obtaining the matrix $A^{\star\star}$ for the set $K^{\prime}=K\cap B_{2}(0,\tilde{d}/2)$ we transform the data to $(A^{\star\star}Y_{i})_{i\in[N]}$ and we apply the polynomial-time algorithm of depersin2022 to the data with $k=u=\lceil N\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge N$ blocks. In addition, when we obtain the matrix $A^{\star\star}$ we use $N^{-1/2}\sigma$ in place of $\sigma$ in Theorem 2.5 and Algorithm 2.
@@
The estimator $f((A^{\star\star}Y_{i})_{i\in[N]})$, where $f$ denotes the procedure of depersin2022 with $k=u=\lceil N\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge N$, of $A^{\star\star}\mu$ satisfies
@@
From $\tilde{d}\geq C\sigma/\sqrt{N}$ we have that $\lceil N\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\asymp N\tilde{d}^{2}/(C^{2}\sigma^{2})$. Since $A^{\star\star}\tilde{Y}=A^{\star\star}\mu+A^{\star\star}\xi$, we will first evaluate $\operatorname{tr}(\operatorname{Var}(A^{\star\star}\xi))$ and $\|\operatorname{Var}(A^{\star\star}\xi)\|_{\operatorname{op}}$ as they appear on the right hand side of the main inequality in Theorem 2 of depersin2022.
@@
This, in conjunction with the result of depersin2022, completes the proof, after the use of the union bound. ∎
Clearly, we can now prove an analogous result to Lemma 2.9 for this setting, with $A_{j}^{\star\star}Y$ substituted with $f((A_{j}^{\star\star}Y_{i})_{i\in[N]})$ with $k=u$ as specified in Lemma 4.2. For the sake of brevity we omit stating this lemma, noting that we denote the analogous constant as $L$. In addition, using Theorem 2 of depersin2022 with $k=u=N$, on the untransformed data $(Y_{i})_{i\in[N]}$, provided that $\varepsilon$ is smaller than an absolute constant $c_{0}$, with probability at least $1-\exp(-N/C_{0})$ we have
@@ 4 Near-optimal Robust Mean Estimation Algorithm
3: Set $\hat{\mu}_{1}:=\Pi_{K}f((Y_{i})_{i\in[N]})$ with $\epsilon=\sigma$ in Lemma A.1 and $k=u=N$ (here $f$ denotes the procedure of depersin2022)
@@
6:  Find $\tilde{\mu}_{j+1}=(f((A_{j}^{\star\star}Y_{i})_{i\in[N]})-A_{j}^{\star\star}\hat{\mu}_{j})/2$ with $k=u=\lceil N\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})\rceil\wedge N$ for some large constant $C>C^{\prime}$ (here $f$ denotes the procedure of depersin2022)
@@
The proof of this result follows the same strategy as the proof of Theorem 2.11, hence we omit it. We now provide a quick sketch of the main changes required. Since here we are assuming $\sigma$ is relatively large, i.e., $\log(1+1/\sigma)$ is polynomial in $n$ and $N$ we can afford to set $r$ to a smaller value in Algorithm 4. The main difference in the two proofs stems from the fact that in the robust case, one needs to assume $k/N=\lceil\tilde{d}_{j}/(C^{2}\sigma^{2})\rceil\wedge 1\gtrsim\epsilon$, but this condition can be handled as in the proof of Theorem 2.11. This is also precisely the reason why we have a maximum with $\sqrt{\varepsilon}\sigma$ in the definition of $\eta^{\star}$. Also, observe that since for our pilot estimator, $\tilde{d}_{1}\lesssim\sigma$, the number of blocks $k$ will be of the order of $N\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})$ almost always, which implies that the high probability guarantees will be of the form $1-\exp(-N\tilde{d}_{j}^{2}/(C^{\prime}\sigma^{2}))-1/(1+N^{3/2}\tilde{d}_{j}^{3}/\sigma^{3})$ for some $C^{\prime}$. The only exception for this is that the probability bound is capped by $1-\exp(-N/C_{0})-1/(1+N^{3/2}\tilde{d}_{j}^{3}/\sigma^{3})$ when $\tilde{d}_{j}\gtrsim\sigma$, by the result of depersin2022. But since our estimator is proper even if $\|\hat{\mu}-\mu\|_{2}^{2}\gtrsim\sigma^{2}$ we always have the bound $\|\hat{\mu}-\mu\|_{2}^{2}\leq d^{2}$. Hence, when we change the high probability bounds part (2.11), and the integration afterwards, we will have an extra term which is of the form $d^{2}\exp(-N/C_{0})$. Thus if $d^{2}\exp(-N/C_{0})\lesssim\sigma^{2}/N\lesssim\eta^{\star 2}$ we can fold that term in the minimax rate, which is why we assumed it in the theorem. It is also worth noting that the minimax rate on the squared level over the ball $B_{2}(0,r)$ (where here we use the value of $r$ before we overwrite it) is $(n/N\sigma^{2}\vee\epsilon\sigma^{2})\wedge r^{2}$, by a result of peng2025noisy. Hence, once we reach $\tilde{d}_{j}\leq 2r$ (with the latter being the overwritten value of $r$) it follows that $\tilde{d}_{j}$ is below the minimax rate of the set $K$.
@@ 5 Linear Regression
Suppose the covariates $Z_{i}$ are centered [^17] If the predictors $Z_{i}$ are not centered one can consider $(Y_{2i}-Y_{2i-1},Z_{2i}-Z_{2i-1})_{i\in[\lfloor N/2\rfloor]}$ to center the predictors while leave the $\beta$ unchanged. We note that this operation prevents the model to have an intercept. sub-Gaussian variables with a well conditioned covariance matrix $\mathbb{E}Z_{i}Z_{i}^{\top}=\Sigma$, i.e., $c^{\prime}\leq\lambda_{\min}(\Sigma)\leq\lambda_{\max}(\Sigma)\leq C^{\prime}$ for some $c^{\prime},C^{\prime}>0$, and sub-Gaussian parameter of constant order, i.e. $\sup_{v\in S^{n-1}}\mathbb{E}\exp(\lambda v^{\top}Z_{i})\leq\exp(\lambda^{2}\zeta^{2}/2)$ for $\zeta=O(1)$. Then the minimax rate $\eta^{\star 2}$, in the case when $\xi_{i}\sim N(0,1)$, is given by the entropy equation $\eta^{\star}=\sup_{\eta}\{\eta\geq 0:N\eta^{2}\leq\log M_{K}^{\operatorname{loc}}(\eta)\}$, as shown in prasadan2025characterizingminimaxratenonparametric.
@@
We consider the estimate $\beta^{\prime}_{j+1}$ defined in (5.1).
We will argue that by Theorem 2.5.9 of dadush2012integer (see also Theorem 1 in lee2018efficient) that $\beta^{\prime}_{j+1}$ is computable. First, we note that the set $K_{(j)}$ admits a membership oracle since we can evaluate its Minkowski gauge. Furthermore, the convex map $\nu\mapsto\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})-Z_{i}^{\top}A^{\star\star}\nu)_{i\in[N]}\|_{2}$ is Lipschitz. We have
@@
We have by Theorem 5.39 of vershynin2010introduction:
@@
$\sqrt{\lambda_{\max}\bigg(\sum_{i\in[N]}Z_{i}Z_{i}^{\top}\bigg)}\lesssim\sqrt{N}+C\sqrt{n}+\sqrt{N}\tilde{d}_{j},$
and thus the map $\nu\mapsto\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})-Z_{i}^{\top}A^{\star\star}\nu)_{i\in[N]}\|_{2}$ is Lipschitz with a constant $\lesssim(\sqrt{N}+c\sqrt{n}+\sqrt{N}\tilde{d}_{j})\asymp(\sqrt{N}+\sqrt{n})$, since $\tilde{d}_{j}=O(1)$ by assumption. Hence the above approximate minimization over $K_{(j)}$ is computable by Theorem 2.5.9 in dadush2012integer.
Let $K^{\prime}_{(j)}=\{A_{j}^{\star\star}\nu:\nu\in K_{(j)}\}$. Denote with $\nu_{j}:=(\beta-\hat{\beta}_{j})/2\in K_{(j)}$ by assumption. Let $\Pi_{K^{\prime}_{(j)}}\nu_{j}=\mathop{\mathrm{argmin}}_{\nu\in K^{\prime}_{(j)}}\|\nu-\nu_{j}\|_{2}$. We know that $\|\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j}\|_{2}\leq\|A_{j}^{\star\star}\nu_{j}-\nu_{j}\|_{2}\lesssim\tilde{d}_{j}/c$ where the last inequality follows from a similar argument as in the proof of Lemma 2.9. Next using Theorem 2.5.9 dadush2012integer with $\epsilon=\sqrt{N}\tilde{d}_{j}/C^{\prime}$ we can obtain a near opitmal solution $\beta^{\prime}_{j+1}$. Then by the (near) optimality we have:
$\sum_{i\in[N]}((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}A_{j}^{\star\star}\beta^{\prime}_{j+1})^{2}\leq 2\sum_{i\in[N]}((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}\Pi_{K^{\prime}_{(j)}}\nu_{j})^{2}+2N\tilde{d}_{j}^{2}/C^{{}^{\prime}2}.$
Rearranging and using the shorthand $\hat{\Sigma}:=N^{-1}\sum_{i\in[N]}Z_{i}Z_{i}^{\top}$ we have
@@
$\leq 2(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})^{\top}\hat{\Sigma}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})+2\bigg|2N^{-1}\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\tilde{\beta}_{j+1})\bigg|+2\tilde{d}_{j}^{2}/C^{{}^{\prime}2},$
@@
$\|\hat{\Sigma}^{1/2}(\tilde{\beta}_{j+1}-\nu_{j})\|_{2}\leq\sqrt{2}\bigg(\|\hat{\Sigma}^{1/2}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})\|_{2}+\sqrt{\bigg|2N^{-1}\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\tilde{\beta}_{j+1})\bigg|}+\tilde{d}_{j}/C^{\prime}\bigg).$ (5.2)
@@
where in the last inequality we used that $\tilde{d}_{j}\geq C/\sqrt{N}$. Since both $\xi_{i}$ and $Z_{i}^{\top}v_{j}$ are sub-Gaussian variables their product is sub-exponential, and the constant is $O(1)$ by assumption. Moreover, clearly, $\mathbb{E}\xi_{i}Z_{i}^{\top}v_{k}=0$. Thus
@@
by Proposition 5.16 of vershynin2010introduction. Set $t=\sqrt{N}\tilde{d}_{j}$, and note that $\tilde{d}_{j}=O(1)$ thus:
@@
Once again by Theorem 5.39 of vershynin2010introduction, we have that

### v2

@@ 4 Robust Heavy-Tailed Extension
We are interested in estimating $\mu$ in a minimax optimal way. In order to derive the rate Peng et al. (2025), suppose that $N\gtrsim\sup_{\delta>0}\log M^{\operatorname{loc}}_{K}(\delta)$. Since $n\gtrsim\log M^{\operatorname{loc}}_{K}(\delta)$ this assumption is implied if $N\gtrsim n$. On the other hand, since we assume that the set $K$ is well-balanced, we have that $rB_{2}\subseteq K$ and therefore, when $\delta\lesssim r$, $\log M^{\operatorname{loc}}_{K}(\delta)\gtrsim n$. Thus, in this section, we will assume the assumption: $N\gtrsim n$. In that case, as shown by Peng et al. (2025) the minimax rate $\inf_{\hat{\mu}}\sup_{\mu\in K}\sup_{\mathcal{C}}\mathbb{E}\|\hat{\mu}-\mu\|_{2}^{2}$ is $\asymp\eta^{\star 2}\wedge d^{2}$ where $\eta^{\star}=(\delta^{\star}\vee\sqrt{\varepsilon}\sigma)$ where $\delta^{\star}:=\sup\{\delta\geq 0:N\delta^{2}/\sigma^{2}\leq\log M^{\operatorname{loc}}_{K}(\delta)\}$, and where the inner most supremum over $\mathcal{C}$ indicates taking supremum over all possible adversarial schemes.
@@
We will modify several steps in our algorithm. Specifically, after obtaining the matrix $A^{\star\star}$ for the set $K^{\prime}=K\cap B_{2}(0,\tilde{d}/2)$ we transform the data to $(A^{\star\star}Y_{i})_{i\in[N]}$ and we apply the polynomial-time algorithm of Depersin and Lecué (2022) to the data with $k=u=\lceil N\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge N$ blocks. In addition, when we obtain the matrix $A^{\star\star}$ we use $N^{-1/2}\sigma$ in place of $\sigma$ in Theorem 2.5 and Algorithm 2.
@@
The estimator $f((A^{\star\star}Y_{i})_{i\in[N]})$, where $f$ denotes the procedure of Depersin and Lecué (2022) with $k=u=\lceil N\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge N$, of $A^{\star\star}\mu$ satisfies
@@
From $\tilde{d}\geq C\sigma/\sqrt{N}$ we have that $\lceil N\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\asymp N\tilde{d}^{2}/(C^{2}\sigma^{2})$. Since $A^{\star\star}\tilde{Y}=A^{\star\star}\mu+A^{\star\star}\xi$, we will first evaluate $\operatorname{tr}(\operatorname{Var}(A^{\star\star}\xi))$ and $\|\operatorname{Var}(A^{\star\star}\xi)\|_{\operatorname{op}}$ as they appear on the right hand side of the main inequality in Theorem 2 of Depersin and Lecué (2022).
@@
This, in conjunction with the result of Depersin and Lecué (2022), completes the proof, after the use of the union bound. ∎
Clearly, we can now prove an analogous result to Lemma 2.9 for this setting, with $A_{j}^{\star\star}Y$ substituted with $f((A_{j}^{\star\star}Y_{i})_{i\in[N]})$ with $k=u$ as specified in Lemma 4.2. For the sake of brevity we omit stating this lemma, noting that we denote the analogous constant as $L$. In addition, using Theorem 2 of Depersin and Lecué (2022) with $k=u=N$, on the untransformed data $(Y_{i})_{i\in[N]}$, provided that $\varepsilon$ is smaller than an absolute constant $c_{0}$, with probability at least $1-\exp(-N/C_{0})$ we have
@@ 4 Near-optimal Robust Mean Estimation Algorithm
3: Set $\hat{\mu}_{1}:=\Pi_{K}f((Y_{i})_{i\in[N]})$ with $\epsilon=\sigma$ in Lemma A.1 and $k=u=N$ (here $f$ denotes the procedure of Depersin and Lecué (2022))
@@
6:  Find $\tilde{\mu}_{j+1}=(f((A_{j}^{\star\star}Y_{i})_{i\in[N]})-A_{j}^{\star\star}\hat{\mu}_{j})/2$ with $k=u=\lceil N\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})\rceil\wedge N$ for some large constant $C>C^{\prime}$ (here $f$ denotes the procedure of Depersin and Lecué (2022))
@@
The proof of this result follows the same strategy as the proof of Theorem 2.11, hence we omit it. We now provide a quick sketch of the main changes required. Since here we are assuming $\sigma$ is relatively large, i.e., $\log(1+1/\sigma)$ is polynomial in $n$ and $N$ we can afford to set $r$ to a smaller value in Algorithm 4. The main difference in the two proofs stems from the fact that in the robust case, one needs to assume $k/N=\lceil\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})\rceil\wedge 1\gtrsim\epsilon$, but this condition can be handled as in the proof of Theorem 2.11. This is also precisely the reason why we have a maximum with $\sqrt{\varepsilon}\sigma$ in the definition of $\eta^{\star}$. Also, observe that since for our pilot estimator, $\tilde{d}_{1}\lesssim\sigma$, the number of blocks $k$ will be of the order of $N\tilde{d}_{j}^{2}/(C^{2}\sigma^{2})$ almost always, which implies that the high probability guarantees will be of the form $1-\exp(-N\tilde{d}_{j}^{2}/(C^{\prime}\sigma^{2}))-1/(1+N^{3/2}\tilde{d}_{j}^{3}/\sigma^{3})$ for some $C^{\prime}$. The only exception for this is that the probability bound is capped by $1-\exp(-N/C_{0})$ when $\tilde{d}_{j}\gtrsim\sigma$, by the result of Depersin and Lecué (2022) (observe that here we do not have the term $1/(1+N^{3/2}\tilde{d}_{j}^{3}/\sigma^{3})$ since this term indicates the failure probability for the matrix $A_{j}^{\star\star}$, and if that happens by the logic of the proof of Theorem 2.11, the estimator will have a rate proportional to that of the pilot estimator which is $\lesssim\sigma$ with probability at least $1-\exp(-N/C_{0})$). But since our estimator is proper even if $\|\hat{\mu}-\mu\|_{2}^{2}\gtrsim\sigma^{2}$ we always have the bound $\|\hat{\mu}-\mu\|_{2}^{2}\leq d^{2}$. Hence, when we change the high probability bounds part (2.11), and the integration afterwards, we will have an extra term which is of the form $d^{2}\exp(-N/C_{0})$. Thus if $d^{2}\exp(-N/C_{0})\lesssim\sigma^{2}/N\lesssim\eta^{\star 2}$ we can fold that term in the minimax rate, which is why we assumed it in the theorem. It is also worth noting that the minimax rate on the squared level over the ball $B_{2}(0,r)$ (where here we use the value of $r$ before we overwrite it) is $(n/N\sigma^{2}\vee\epsilon\sigma^{2})\wedge r^{2}$, by a result of Peng et al. (2025). Hence, once we reach $\tilde{d}_{j}\leq 2r$ (with the latter being the overwritten value of $r$) it follows that $\tilde{d}_{j}$ is below the minimax rate of the set $K$.
@@ 5 Linear Regression
Suppose the covariates $Z_{i}$ are centered [^17] If the predictors $Z_{i}$ are not centered one can consider $(Y_{2i}-Y_{2i-1},Z_{2i}-Z_{2i-1})_{i\in[\lfloor N/2\rfloor]}$ to center the predictors while leave the $\beta$ unchanged. We note that this operation prevents the model from having an intercept. sub-Gaussian variables with a well conditioned covariance matrix $\mathbb{E}Z_{i}Z_{i}^{\top}=\Sigma$, i.e., $c^{\prime}\leq\lambda_{\min}(\Sigma)\leq\lambda_{\max}(\Sigma)\leq C^{\prime}$ for some $c^{\prime},C^{\prime}>0$, and sub-Gaussian parameter of constant order, i.e. $\sup_{v\in S^{n-1}}\mathbb{E}\exp(\lambda v^{\top}Z_{i})\leq\exp(\lambda^{2}\zeta^{2}/2)$ for $\zeta=O(1)$. Then the minimax rate $\eta^{\star 2}$, in the case when $\xi_{i}\sim N(0,1)$, is given by the entropy equation $\eta^{\star}=\sup_{\eta}\{\eta\geq 0:N\eta^{2}\leq\log M_{K}^{\operatorname{loc}}(\eta)\}$, as shown in Prasadan and Neykov (2025a).
@@
We consider the estimate $\beta^{\prime}_{j+1}$ defined in (5.1). We will argue that by Theorem 2.5.9 of Dadush (2012) (see also Theorem 1 in Lee et al. (2018)) that $\beta^{\prime}_{j+1}$ is computable. First, we note that the set $K_{(j)}$ admits a membership oracle since we can evaluate its Minkowski gauge. Furthermore, the convex map $\nu\mapsto\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})-Z_{i}^{\top}A^{\star\star}\nu)_{i\in[N]}\|_{2}$ is Lipschitz. We have
@@
We have by Theorem 5.39 of Vershynin (2010):
@@
$\sqrt{\lambda_{\max}\bigg(\sum_{i\in[N]}Z_{i}Z_{i}^{\top}\bigg)}\lesssim\sqrt{N}+C\sqrt{n}+\sqrt{N}\tilde{d}_{j},$ (5.2)
and thus the map $\nu\mapsto\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})-Z_{i}^{\top}A^{\star\star}\nu)_{i\in[N]}\|_{2}$ is Lipschitz with a constant $\lesssim(\sqrt{N}+c\sqrt{n}+\sqrt{N}\tilde{d}_{j})\asymp(\sqrt{N}+\sqrt{n})$, since $\tilde{d}_{j}=O(1)$ by assumption. Hence the above approximate minimization over $K_{(j)}$ is computable by Theorem 2.5.9 in Dadush (2012).
Let $K^{\prime}_{(j)}=\{A_{j}^{\star\star}\nu:\nu\in K_{(j)}\}$. Denote with $\nu_{j}:=(\beta-\hat{\beta}_{j})/2\in K_{(j)}$ by assumption. Let $\Pi_{K^{\prime}_{(j)}}\nu_{j}=\mathop{\mathrm{argmin}}_{\nu\in K^{\prime}_{(j)}}\|\nu-\nu_{j}\|_{2}$. We know that $\|\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j}\|_{2}\leq\|A_{j}^{\star\star}\nu_{j}-\nu_{j}\|_{2}\lesssim\tilde{d}_{j}/c$ where the last inequality follows from a similar argument as in the proof of Lemma 2.9. Next using Theorem 2.5.9 Dadush (2012) with $\epsilon\asymp\sqrt{N}\tilde{d}_{j}^{2}$ we can obtain a near opitmal solution $\beta^{\prime}_{j+1}$. Then by the (near) optimality we have:
$\sum_{i\in[N]}((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}A_{j}^{\star\star}\beta^{\prime}_{j+1})^{2}$ $\leq\sum_{i\in[N]}((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}\Pi_{K^{\prime}_{(j)}}\nu_{j})^{2}+\epsilon^{2}$
$+2\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}\Pi_{K^{\prime}_{(j)}}\nu_{j})_{i\in[N]}\|_{2}\epsilon.$ (5.3)
Next using the shorthand $\hat{\Sigma}:=N^{-1}\sum_{i\in[N]}Z_{i}Z_{i}^{\top}$, we handle the last term of the right hand side:
$\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}\Pi_{K^{\prime}_{(j)}}\nu_{j})_{i\in[N]}\|_{2}\leq 2\sqrt{N(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})^{\top}\hat{\Sigma}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})}+2\|\xi\|_{2}.$
Using (5.2), the first term can be bounded as $\sqrt{N(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})^{\top}\hat{\Sigma}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})}\lesssim(\sqrt{N}+\sqrt{n})\tilde{d}_{j}/c$. On the other hand, by Proposition 5.16 of Vershynin (2010) we know that
$\mathbb{P}(\|\xi\|^{2}_{2}-\mathbb{E}\|\xi\|^{2}_{2}\geq t)\leq 2\exp(-C^{\prime\prime}(t^{2}/N\wedge t)).$
Hence we can select $t=N\tilde{d}_{j}$ to obtain that with probability at least $1-2\exp(-C^{\prime\prime}N\tilde{d}_{j}^{2})$ we have $\|\xi\|^{2}_{2}\leq\mathbb{E}\|\xi\|_{2}^{2}+N\tilde{d}_{j}\leq N\gamma^{2}+N\tilde{d}_{j}\lesssim N$. Hence since $\epsilon\asymp\sqrt{N}\tilde{d}_{j}^{2}$, and $\tilde{d}_{j}=O(1)$ we have that:
$\epsilon^{2}+2\|((Y_{i}-Z_{i}^{\top}\hat{\beta}_{j})/2-Z_{i}^{\top}\Pi_{K^{\prime}_{(j)}}\nu_{j})_{i\in[N]}\|_{2}\epsilon=O(N\tilde{d}_{j}^{2}).$
with high probability.
Then by rearranging (5) we have
@@
$\leq(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})^{\top}\hat{\Sigma}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})+2\bigg|N^{-1}\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\tilde{\beta}_{j+1})\bigg|+\tilde{d}_{j}^{2}/C^{{}^{\prime}2},$
@@
$\|\hat{\Sigma}^{1/2}(\tilde{\beta}_{j+1}-\nu_{j})\|_{2}\leq\|\hat{\Sigma}^{1/2}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\nu_{j})\|_{2}+\sqrt{\bigg|2N^{-1}\sum_{i\in[N]}\xi_{i}Z_{i}^{\top}(\Pi_{K^{\prime}_{(j)}}\nu_{j}-\tilde{\beta}_{j+1})\bigg|}+\tilde{d}_{j}/C^{\prime}.$ (5.4)
@@
where in the last inequality we used that $\tilde{d}_{j}\geq C/\sqrt{N}$. Since both $\xi_{i}$ and $Z_{i}^{\top}v_{k}$ are sub-Gaussian variables their product is sub-exponential, and the constant is $O(1)$ by assumption. Moreover, clearly, $\mathbb{E}\xi_{i}Z_{i}^{\top}v_{k}=0$. Thus
@@
by Proposition 5.16 of Vershynin (2010). Set $t=\sqrt{N}\tilde{d}_{j}$, and note that $\tilde{d}_{j}=O(1)$ thus:
@@
Once again by Theorem 5.39 of Vershynin (2010), we have that